Vertex form of y = 4x^2+12x+8. Please show your solution.
Answer by htmentor(1343)

Vertex form is y = a(x-h)^2 + k, where (h,k) is the vertex
The x-coordinate of the vertex, h = -b/2a = -12/8 = -3/2
Thus, k = 4(-3/2)^2 + 12(-3/2) + 8 = 4(9/4) - 10 = -1
So we have y = 4(x+3/2)^2 - 1